### v1

@@ 2.2 Deep LSTM Network
Each LSTM layer consists of a set of LSTM cells, as depicted in Figure 3. Like the neural nodes in traditional ANNs, each LSTM cell contains an independent set of weights and biases shared across the entire temporal sequence within the layer. First, the basic architecture of the deep LSTM network for sequence-to-sequence modeling is presented[zhang2019deep]. The architecture includes a number of hidden layers - additional to input and output layers - as shown in Figure 3(a), some of which are LSTM layers, while others are fully connected layers. The LSTM cell includes four components: an internal cell, an input gate, a forget gate, and an output gate. The internal cell remembers the cell state from the previous time step via a self-recurrent connection. The input gate maintains the flow of input activations into the internal cell state, while the output gate controls the flow of output activations from the LSTM cell. The forget gate scales the internal cell state, allowing the LSTM cell to adaptively forget or reset its memory. Through the coordinated actions of the input, forget, and output gates, the cell state can selectively transmit essential information along the temporal sequence, capturing both long-term and short-term dependencies in a dynamic system effectively.
@@
Where $W_{(\alpha\beta)}^{(l)}$ (with $\alpha$ = $\{x,h\}$} and $\beta={\{f,i,c,o\}}$ ) denotes weight matrix corresponding to different inputs (e.g. $x_{t}^{l}$ or $h_{t}^{l}$ ) different gates (e.g., input gate, forget gate, tanh layer, or output gate). And $b_{\beta}^{(l)}$ represents the corresponding bias vectors; the lth layer of the LSTM network is denoted by the superscript l. Here, $\tilde{c}_{t}^{(l)}$ denotes a vector of intermediate candidate values created by a tanh layer shown in Fig. 1(b); $\sigma$ is the logistic sigmoid function; tanh is the hyperbolic tangent function; $\odot$ denotes the Hadamard product (elementwise product). The complex connection system within each LSTM cell makes the deep LSTM network powerful in sequence modelling, in which the fully connected layers are beneficial to map the temporal feature maps to the corresponding output space.
The output of the final dense layer produces a time-series prediction of the state-space variables $z(t)={x(t),\dot{x}(t),g(t)}$, where $x$ is displacement, $\dot{x}$ is velocity, and $g$ is restoring force. While this formulation accurately captures temporal patterns, it lacks explicit knowledge of the governing physical laws.To address this limitation, the predicted outputs are further processed using a graph-based tensor differentiator to compute their temporal derivatives. A specialized loss function is then formulated, tailored to the specific requirements of each scenario. The tensor differentiator and the loss function are described in detail in Sections 2.3 and  2.4, respectively. Training proceeds by minimizing the loss:
@@
The model is implemented in Python using Keras [chollet2018keras] with TensorFlow backend. Training is conducted on Google Colab Pro with GPU acceleration. All hyperparameters, including the number of LSTM layers $m$, number of hidden units per layer, batch size, learning rate, and optimizer settings, are specified in the Supplementary Information.The detailed network architecture of the proposed framework is represented in Figure 1.
2.3 Graph Based Tensor Differentiator
Numerical differentiation is an essential method widely used in computational mathematics, particularly in solving differential equations and gradient-based optimization. Finite differences (FD) approximate derivatives by combining nearby function values using a set of weights. Here in this study to calculate the derivatives of the state space outputs of the neural network a graph-based tensor differentiator utilizing finite difference approach is implemented.[phycnn] This section gives detailed explanation about the construction of finite difference matrix ($\Phi_{t}]$) to calculate the first order derivatives providing second order accuracy meaning the truncation error is of order 2 (i.e. $O(\Delta t^{2})$). For all the interior points central difference stencil is utilized whereas for the first and last point forward and backward difference stencil is implemented.
@@ 2.4 Loss function
In this study, three different loss functions are utilised for each of the two numerical validation scenario and the experimental validation of the model. To incorporate physical constraints, temporal derivatives $\dot{z}(t)={x_{t}(t),\dot{x}_{t}(t),g_{t}(t)}$ are estimated using a graph-based differentiator [zhang2019deep], which implements finite-difference approximations over a local stencil. This component enables embedding of physics-based residuals into the loss function. For the first scenario of the numerical validation where full-state measurements are available for training, the total loss function used for training combines data loss and physics-based loss:
$J(\theta)=w_{1}\cdot J_{D}(\theta)+w_{2}\cdot J_{P}(\theta)$ (12)
$J_{D}(\theta)=\frac{1}{N}\left\|x^{p}-x^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}^{p}-\dot{x}^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|g^{p}-g^{m}\right\|_{2}^{2}$ (13)
$J_{P}(\theta)=\frac{1}{N}\left\|\dot{x}^{p}-x_{t}^{p}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}+g^{p}+\Gamma\ddot{x}_{g}\right\|_{2}^{2}$ (14)
@@
Where $J_{D}(\theta)$ is the data loss based on the measurements and $J_{P}(\theta)$ denotes the physics loss . Superscripts $p$ and $m$ denote predicted and measured quantities, respectively. When certain measurement channels are unavailable, corresponding loss terms are omitted. In the physics loss the second term incorporates the physical laws in terms of the governing equation where $\dot{x}_{t}^{p}$ represents the derivative of the predicted velocity (which corresponds to the acceleration). The first term serves as a consistency or sanity check, ensuring that the predicted state-space variables align with their physical definitions. Specifically, the predicted velocity $\dot{x}^{p}$ should be equal with the derivative of the predicted displacement($x_{t}^{p}$) computed using the graph-based tensor differentiator. In this study, equal weights ($w_{1}=w_{2}=1$) are used, but these can be tuned depending on the application. For the second case of numerical validation, although the system information—such as the mass matrix, stiffness, and damping coefficients—is available, only acceleration measurements are used as training data. A modified loss function is architected to incorporate the governing equations and enable training using acceleration data alone to accommodate this limitation. The formulation of this loss function is presented below:
$J(\theta)=\frac{1}{N}\left\|\dot{x}^{p}-x_{t}^{p}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}-\ddot{x}^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}+g^{p}+\Gamma\ddot{x}_{g}\right\|_{2}^{2},$ (15)
This modified loss function’s first term is identical to the original physics loss’s first term, ensuring consistency between predicted state variables. The third term also remains unchanged from the second term of the original physics loss, enforcing adherence to the physical laws. In the second term the $\dot{x}_{t}^{p}$ represents the time derivative of the predicted velocity, computed via the graph-based tensor differentiator, and enforced to match the available measured acceleration data.
@@
For the experimental validation where we dont have a complete description of the structural system—including mass, stiffness, damping properties, and interaction effect, it is not possible to implement physics based loss term and the model relies solely on the data loss. In this purely data-driven scenario, only acceleration data gathered from the sensors are available for training. The model is designed to predict displacements, which are then differentiated twice using the graph-based tensor differentiator to obtain the corresponding accelerations ($x_{tt}$) using tensor differentiator. The loss function implemented for this case is presented below:
$J(\theta)=\frac{1}{N}\left\|\ddot{x}^{m}-x_{tt}^{p}\right\|_{2}^{2}$ (16)
2.5 Numerical Validation:
@@
where $m=1\,\mathrm{kg}$, $c=1\,\mathrm{Ns/m}$, $k_{1}=20\,\mathrm{N/m}$, and $k_{2}=200\,\mathrm{N/m}$. The restoring force $h$ is normalized by mass to yield $g=h/m$, following Zhang [phycnn]. Two scenarios are considered. In the first, full-state measurements—displacement, velocity, and restoring force—are assumed available for training. These can be extracted from numerical simulations, with $g$ computed via $g=-\ddot{x}-\Gamma\ddot{x}_{g}$. In the second, only acceleration data are used, representing a more constrained condition. This setup enables direct assessment of the proposed model’s robustness under varying levels of data availability.
2.5.1 Case 1: Full-state measurements available for training
We consider the scenario in which full-state measurements—displacement ($x$), velocity ($\dot{x}$), and restoring force (gg)—are available for training the model. Following Zhang [phycnn], we adopt a dataset of 100 numerical simulations of a nonlinear single-degree-of-freedom system subjected to synthetic earthquake ground motions drawn from the PEER Strong Motion Database [chiou2008nga], representing a 10% probability of exceedance in 50 years. Each simulation spans 50 seconds with a sampling rate of 20 Hz, resulting in 1001 time steps per example. Out of these, 10 records are randomly selected for training, while the remaining 90 are reserved for validation. The proposed PhyULSTM framework (Fig. 1) is trained to map ground motion acceleration to system response using these data. The model input and output tensors are of shapes [10, 1001, 1] and [10, 1001, 3], respectively. Training follows the architecture and loss function detailed in 2.4.
@@ 2.5.1 Case 1: Full-state measurements available for training
$J_{D}(\theta)=\frac{1}{N}\left\|x^{p}-x^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}^{p}-\dot{x}^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|g^{p}-g^{m}\right\|_{2}^{2}$
$J_{P}(\theta)=\frac{1}{N}\left\|\dot{x}^{p}-x_{t}^{p}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}+g^{p}+\Gamma\ddot{x}_{g}\right\|_{2}^{2}$
@@
Figure 4 illustrates the predicted displacement response for three unseen ground motion using both PhyULSTM and PhyCNN models. The proposed PhyULSTM architecture demonstrates excellent predictive performance, accurately reproducing the displacement time history with close agreement to the ground truth. Notably, PhyULSTM captures both the overall trend and the rapid fluctuations in the response that are significantly underrepresented by PhyCNN. This difference is especially prominent in regions with sharp nonlinear behavior, underscoring PhyULSTM’s superior ability to generalize across previously unseen inputs. Figure 5 further compares the predicted velocity and nonlinear restoring force for the a representative ground motion. Again, PhyULSTM closely matches the true values obtained from direct numerical integration of the governing equations. In contrast, PhyCNN tends to smooth out higher-frequency variations and fails to capture abrupt transitions in the restoring force dynamics. This limitation becomes particularly evident in the high-gradient regions, where PhyULSTM maintains fidelity to the true signal.
@@
The nonlinear dynamic behavior of the system is further examined through hysteresis plots of the restoring force versus displacement and velocity, as shown in Fig. 6. These plots serve as critical indicators of the model’s ability to capture complex energy dissipation mechanisms and stiffness degradation inherent in nonlinear structural systems.
@@
Figure 4: Time histories of displacement responses predicted by PhyULSTM (red, left column) and PhyCNN (red, right column) compared against the reference solution (blue). Each row corresponds to a different earthquake input (1, 2, and 3). Subplots are labeled 1a, 1b, 2a, 2b, 3a, and 3b for reference, where the left subplots (a) show PhyULSTM predictions and the right subplots (b) show PhyCNN predictions. Across all three test cases, PhyULSTM consistently tracks the true system behavior with higher accuracy, capturing both transient peaks and steady-state oscillations. In contrast, PhyCNN fails to reproduce finer-scale dynamics and underestimates displacement amplitudes during critical phases.
The results in Fig. 6 thus confirm that PhyULSTM is not only capable of accurate time-series predictions but also preserves the nonlinear physics of the underlying system, making it a reliable tool for surrogate modeling in structural dynamics.
@@
Figure 5: Predicted time histories of velocity ($\dot{x}$) and nonlinear restoring force ($g$) for a representative ground motion are shown.The left column shows the responses predicted by PhyULSTM, and the right column shows those predicted by PhyCNN. Subplots are labeled 4a, 4b, 4c, and 4d for reference: 4a and 4c correspond to velocity and normalized restoring force predicted by PhyULSTM, while 4b and 4d correspond to the same predicted by PhyCNN. The PhyULSTM model captures both quantities with high fidelity compared to the numerical reference, successfully reproducing subtle fluctuations and transient peaks.
@@
Figure 7 presents the results of a regression analysis comparing the predicted and true displacement responses for both PhyULSTM and PhyCNN models. The proposed PhyULSTM framework demonstrates remarkable predictive fidelity, achieving a maximum correlation coefficient of 0.998 and a minimum of 0.83 across all 90 unseen test cases. These values indicate near-perfect agreement in most scenarios, even under conditions involving significant nonlinearity and variability in input ground motions. In contrast, the PhyCNN model exhibits noticeably lower predictive accuracy, with correlation coefficients ranging from 0.97 down to 0.577, suggesting poorer generalization and a tendency to underperform in more complex test cases.
@@
2.5.2 Case 2: Training with acceleration measurements only
Figure 8: Time histories of displacement responses predicted (red) by PhyULSTM (left) and PhyCNN (right) are compared against reference numerical solutions (blue) obtained via direct integration. Each row corresponds to a different earthquake input (1, 2, and 3). Subplots are labeled 1a, 1b, 2a, 2b, 3a, and 3b for reference, where the left subplots (a) show PhyULSTM predictions and the right subplots (b) show PhyCNN predictions. PhyULSTM consistently captures transient dynamics and amplitude variations with high fidelity across three distinct seismic events, whereas PhyCNN exhibits noticeable deviations, especially in peak displacement magnitudes and oscillation patterns.

### v2

@@ 2.2 Deep LSTM Network
Each LSTM layer consists of a set of LSTM cells, as depicted in Figure 3. Like the neural nodes in traditional ANNs, each LSTM cell contains an independent set of weights and biases shared across the entire temporal sequence within the layer. First, the basic architecture of the deep LSTM network for sequence-to-sequence modeling was presented by [52]. The architecture includes a number of hidden layers - additional to input and output layers - as shown in Figure 3(a), some of which are LSTM layers, while others are fully connected layers. The LSTM cell includes four components: an internal cell, an input gate, a forget gate, and an output gate. The internal cell remembers the cell state from the previous time step via a self-recurrent connection. The input gate maintains the flow of input activations into the internal cell state, while the output gate controls the flow of output activations from the LSTM cell. The forget gate scales the internal cell state, allowing the LSTM cell to adaptively forget or reset its memory. Through the coordinated actions of the input, forget, and output gates, the cell state can selectively transmit essential information along the temporal sequence, capturing both long-term and short-term dependencies in a dynamic system effectively.
@@
Where $W_{(\alpha\beta)}^{(l)}$ (with $\alpha$ = $\{x,h\}$} and $\beta={\{f,i,c,o\}}$ ) denotes weight matrix corresponding to different inputs (e.g. $x_{t}^{l}$ or $h_{t}^{l}$ ) different gates (e.g., input gate, forget gate, tanh layer, or output gate). And $b_{\beta}^{(l)}$ represents the corresponding bias vectors; the lth layer of the LSTM network is denoted by the superscript l. Here, $\tilde{c}_{t}^{(l)}$ denotes a vector of intermediate candidate values created by a tanh layer shown in Figure3(b); $\sigma$ is the logistic sigmoid function; tanh is the hyperbolic tangent function; $\odot$ denotes the Hadamard product (elementwise product). The complex connection system within each LSTM cell makes the deep LSTM network powerful in sequence modelling, in which the fully connected layers are beneficial to map the temporal feature maps to the corresponding output space.
The output of the final dense layer produces a time-series prediction of the state-space variables $z(t)={x(t),\dot{x}(t),g(t)}$, where $x$ is displacement, $\dot{x}$ is velocity, and $g$ is restoring force. While this formulation accurately captures temporal patterns, it lacks explicit knowledge of the governing physical laws.To address this limitation, the predicted outputs are further processed using a tensor differentiator to compute their temporal derivatives. A specialized loss function is then formulated, tailored to the specific requirements of each scenario. The tensor differentiator and the loss function are described in detail in Sections 2.3 and  2.4, respectively. Training proceeds by minimizing the loss:
@@
The model is implemented in Python using Keras [7] with TensorFlow backend. Model training is carried out using GPU acceleration. The detailed network architecture of the proposed framework is represented in Figure 1.
2.3 Tensor Differentiator
Numerical differentiation is an essential method widely used in computational mathematics, particularly in solving differential equations and gradient-based optimization. Finite differences (FD) approximate derivatives by combining nearby function values using a set of weights. Here in this study to calculate the derivatives of the state space outputs of the neural network, a tensor differentiator utilizing finite difference approach is implemented.[53] This section gives detailed explanation about the construction of finite difference matrix ($\Phi_{t}]$) to calculate the first order derivatives providing second order accuracy meaning the truncation error is of order 2 (i.e. $O(\Delta t^{2})$). For all the interior points central difference stencil is utilized whereas for the first and last point forward and backward difference stencil is implemented.
@@ 2.4 Loss function
In this study, three different loss functions are utilised for each of the two numerical validation scenario and the experimental validation of the model. To incorporate physical constraints, temporal derivatives $\dot{z}(t)={x_{t}(t),\dot{x}_{t}(t),g_{t}(t)}$ are estimated using a tensor differentiator [52], which implements finite-difference approximations over a local stencil. This component enables embedding of physics-based residuals into the loss function. For the first scenario of the numerical validation where full-state measurements are available for training, the total loss function used for training combines data loss and physics-based loss:
$J(\theta)=w_{1}J_{D}(\theta)+w_{2}J_{P}(\theta)$ (12)
$J_{D}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\|x^{p(i)}-x^{m(i)}\|_{2}^{2}+\frac{1}{N}\sum_{i=1}^{N}\|\dot{x}^{p(i)}-\dot{x}^{m(i)}\|_{2}^{2}+\frac{1}{N}\sum_{i=1}^{N}\|g^{p(i)}-g^{m(i)}\|_{2}^{2}$ (13)
$J_{P}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{x}^{p(i)}-x_{t}^{p(i)}\right\|_{2}^{2}+\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{x}_{t}^{p(i)}+g^{p(i)}+\Gamma\ddot{x}^{(i)}_{g}\right\|_{2}^{2}$ (14)
@@
Where $J_{D}(\theta)$ is the data loss based on the measurements and $J_{P}(\theta)$ denotes the physics loss. The superscripts $p$ and $m$ denote the predicted and measured (actual) responses, respectively. The index $i$ represents the $i^{\text{th}}$ earthquake sample, and $N$ is the total number of earthquake records used for training. The terms $x$, $\dot{x}$, and $g$ correspond to the structural displacement, velocity, and normalized restoring force, respectively. When certain measurement channels are unavailable, corresponding loss terms are omitted. In the physics loss the second term incorporates the physical laws in terms of the governing equation where $\dot{x}_{t}^{p}$ represents the derivative of the predicted velocity (which corresponds to the acceleration). The first term serves as a consistency or sanity check, ensuring that the predicted state-space variables align with their physical definitions. Specifically, the predicted velocity $\dot{x}^{p}$ should be equal with the derivative of the predicted displacement($x_{t}^{p}$) computed using the tensor differentiator. In this study, equal weights ($w_{1}=w_{2}=1$) are used, but these can be tuned depending on the application. In the other case, although the system information—such as the mass matrix, stiffness, and damping coefficients—is available, only acceleration measurements are used as training data. A modified loss function is architected to incorporate the governing equations and enable training using acceleration data alone to accommodate this limitation. The formulation of this loss function is presented below:
$J_{P}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{x}^{p(i)}-x_{t}^{p(i)}\right\|_{2}^{2}+\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{x}_{t}^{p(i)}+g^{p(i)}+\Gamma\ddot{x}^{(i)}_{g}\right\|_{2}^{2},$ (15)
and
$J_{D}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{x}_{t}^{p(i)}-\ddot{x}^{m(i)}\right\|_{2}^{2}$ (16)
In the data loss, the $\dot{x}_{t}^{p}$ represents the time derivative of the predicted velocity, computed using the tensor differentiator, and enforced to match the available measured acceleration data.
@@
For the experimental validation where we don’t have a complete description of the structural system—including mass, stiffness, damping properties, and interaction effect, it is not possible to implement a physics-based loss term, and the model relies solely on the data loss. In this purely data-driven scenario, only acceleration data gathered from the sensors are available for training. The model is designed to predict displacements, which are then differentiated twice using the tensor differentiator to obtain the corresponding accelerations ($x_{tt}$) using the tensor differentiator. The loss function implemented for this case is presented below:
$J_{D}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left\|x_{tt}^{p(i)}-\ddot{x}^{m(i)}\right\|_{2}^{2}$ (17)
3 Application in Response Prediction of a nonlinear single-degree-of-freedom system:
@@
where $m=1\,\mathrm{kg}$, $c=1\,\mathrm{Ns/m}$, $k_{1}=20\,\mathrm{N/m}$, and $k_{2}=200\,\mathrm{N/m}$. The restoring force $h$ is normalized by mass to yield $g=h/m$, following Zhang [53]. Two scenarios are considered. In the first, full-state measurements—displacement, velocity, and restoring force—are assumed available for training. These can be extracted from numerical simulations, with $g$ computed via $g=-\ddot{x}-\Gamma\ddot{x}_{g}$. In the second, only acceleration data are used, representing a more constrained condition. This setup enables direct assessment of the proposed model’s robustness under varying levels of data availability. The benchmark dataset used for training and validation was obtained from the publicly available PhyCNN repository: https://github.com/zhry10/PhyCNN.
3.1 Case 1: Full-state measurements available for training
We consider the scenario in which full-state measurements—displacement ($x$), velocity ($\dot{x}$), and restoring force (gg)—are available for training the model. Following Zhang [53], we adopt a dataset of 100 numerical simulations of the nonlinear single-degree-of-freedom system18 subjected to synthetic earthquake ground motions drawn from the PEER Strong Motion Database [6], representing a 10% probability of exceedance in 50 years. Each simulation spans 50 seconds with a sampling rate of 20 Hz, resulting in 1001 time steps per example. Out of these, 10 records are randomly selected for training, while the remaining 90 are reserved for validation. The proposed PhyULSTM framework (Figure 1) is trained to map ground motion acceleration to system response using these data. The model input and output tensors are of shapes [10, 1001, 1] and [10, 1001, 3], respectively. Training follows the architecture and loss function detailed in 2.4.
@@ 3.1 Case 1: Full-state measurements available for training
$J_{D}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\|x^{p(i)}-x^{m(i)}\|_{2}^{2}+\frac{1}{N}\sum_{i=1}^{N}\|\dot{x}^{p(i)}-\dot{x}^{m(i)}\|_{2}^{2}+\frac{1}{N}\sum_{i=1}^{N}\|g^{p(i)}-g^{m(i)}\|_{2}^{2}$
$J_{P}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{x}^{p(i)}-x_{t}^{p(i)}\right\|_{2}^{2}+\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{x}_{t}^{p(i)}+g^{p(i)}+\Gamma\ddot{x}_{g}^{(i)}\right\|_{2}^{2}$
@@
To evaluate the generalization capability of the proposed framework under unseen seismic excitations, the model predictions are compared against those obtained from the trained PhyCNN architecture. Figure 4 illustrates the predicted displacement response for three unseen ground motion using both PhyULSTM and PhyCNN models. The proposed PhyULSTM architecture demonstrates excellent predictive performance, accurately reproducing the displacement time history with close agreement to the ground truth. Notably, PhyULSTM captures both the overall trend and the rapid fluctuations in the response that are significantly underrepresented by PhyCNN. This difference is especially prominent in regions with sharp nonlinear behavior, underscoring PhyULSTM’s superior ability to generalize across previously unseen inputs. Figure 5 further compares the predicted velocity and nonlinear restoring force for the a representative ground motion. Again, PhyULSTM closely matches the true values obtained from direct numerical integration of the governing equations. In contrast, PhyCNN tends to smooth out higher-frequency variations and fails to capture abrupt transitions in the restoring force dynamics. This limitation becomes particularly evident in the high-gradient regions, where PhyULSTM maintains fidelity to the true signal.
@@
The nonlinear dynamic behavior of the system is further examined through hysteresis plots of the restoring force versus displacement and velocity, as shown in Figure 6. These plots serve as critical indicators of the model’s ability to capture complex energy dissipation mechanisms and stiffness degradation inherent in nonlinear structural systems. The results in Figure 6 thus confirm that PhyULSTM is not only capable of accurate time-series predictions but also preserves the nonlinear physics of the underlying system, making it a reliable tool for surrogate modeling in structural dynamics.
@@
Figure 4: Time histories of displacement responses predicted by PhyULSTM (red, left column) and PhyCNN (red, right column) compared against the reference solution (blue). Each row corresponds to a different earthquake input (1, 2, and 3). Subplots are labeled 1a, 1b, 2a, 2b, 3a, and 3b for reference, where the left subplots (a) show PhyULSTM predictions and the right subplots (b) show PhyCNN predictions.
@@
Figure 5: Predicted time histories of velocity ($\dot{x}$) and nonlinear restoring force ($g$) for a representative ground motion are shown.The left column shows the responses predicted by PhyULSTM, and the right column shows those predicted by PhyCNN. Subplots are labeled 4a, 4b, 4c, and 4d for reference: 4a and 4c correspond to velocity and normalized restoring force predicted by PhyULSTM, while 4b and 4d correspond to the same predicted by PhyCNN.
@@
Figure 7 presents the results of a regression analysis comparing the predicted and true displacement responses for both PhyULSTM and PhyCNN models. The proposed PhyULSTM framework demonstrates remarkable predictive fidelity, achieving a maximum correlation coefficient of 0.998 and a minimum of 0.83 across all 90 unseen test cases. These values indicate near-perfect agreement in most scenarios, even under conditions involving significant nonlinearity and variability in input ground motions. In contrast, the PhyCNN model exhibits noticeably lower predictive accuracy, with correlation coefficients ranging from 0.97 down to 0.577.
@@
3.1.1 Global error comparison
The overall prediction accuracy for the test set is summarized in Table 2. The PhyULSTM model achieves a mean RMSE of $9.52\times 10^{-3}$, whereas PhyCNN produces a significantly larger value of $2.46\times 10^{-2}$, corresponding to an error reduction of approximately $61\%$. A similar trend is observed for the relative $L_{2}$ error, where PhyULSTM attains $19.28\%$ compared to $47.40\%$ for PhyCNN. The median values follow the same pattern, confirming that the performance improvement is consistent across the dataset rather than being influenced by a small number of samples. These results demonstrate the superior time-domain prediction accuracy of the proposed PhyULSTM framework.
Table 2: Global error metrics for Case 1 (full-state training).
Model Mean RMSE Median RMSE Mean Rel-$L_{2}$ Median Rel-$L_{2}$
PhyULSTM $9.5160\times 10^{-3}$ $5.6856\times 10^{-3}$ $19.28\%$ $15.60\%$
PhyCNN $2.4625\times 10^{-2}$ $1.6814\times 10^{-2}$ $47.40\%$ $41.77\%$
3.1.2 Frequency-domain comparison
To complement the time-domain analysis, Fourier Transform (FFT) analysis was performed on the model predictions and reference responses for 8 random test cases (Fig. 8). The frequency spectra indicate that the system dynamics are dominated by low-frequency components (<2<2 Hz), characterised by pronounced peaks.
Both PhyULSTM and PhyCNN successfully reproduce the dominant frequency locations and the overall spectral decay, indicating that the primary dynamic characteristics of the structure are captured. However, a closer inspection reveals differences in the prediction of peak spectral amplitudes. In several cases, the PhyCNN model either overestimates or underestimates the peak amplitude, whereas the PhyULSTM predictions consistently show closer agreement with the reference spectra.
Figure 8: Comparison of Fourier amplitude spectra of reference and predicted responses for multiple test samples under full-state training.
Accurate estimation of peak response is vital for reliability and fragility analysis, where extreme response quantities govern structural safety. The improved peak amplitude prediction observed for PhyULSTM therefore provides additional evidence of its superior ability to capture structural dynamic behaviour. This frequency-domain validation complements the time-domain error metrics and further confirms the fidelity of the proposed model.
3.1.3 Peak response error analysis
The distribution of peak response errors is illustrated in Fig. 9. The PhyULSTM model achieves a mean peak error of $8.92\%$ and a median error of $5.76\%$, whereas PhyCNN yields significantly larger values of $14.94\%$ and $10.59\%$, respectively. In addition to the lower central tendency, the PhyULSTM results exhibit a much tighter distribution with fewer extreme outliers.
Figure 9: Distribution of peak response errors for the test set under full-state training.
In contrast, the PhyCNN predictions show a wider spread and several large outliers, with peak errors exceeding $100\%$ in some test cases. This behavior indicates reduced robustness in predicting extreme structural responses. The improved peak response accuracy of PhyULSTM is particularly important for structural safety and performance assessment, where accurate estimation of maximum responses is critical.
3.2 Case 2: Training with acceleration measurements only
Figure 10: Time histories of displacement responses predicted (red) by PhyULSTM (left) and PhyCNN (right) are compared against reference numerical solutions (blue) obtained via direct integration. Each row corresponds to a different earthquake input (1, 2, and 3). Subplots are labeled 1a, 1b, 2a, 2b, 3a, and 3b for reference, where the left subplots (a) show PhyULSTM predictions and the right subplots (b) show PhyCNN predictions.